# ECG Semantic Integrator (ESI): A Foundation ECG Model Pretrained with LLM-Enhanced Cardiological Text

### v1

@@ Abstract
The utilization of deep learning on electrocardiogram (ECG) analysis has brought the advanced accuracy and efficiency of cardiac healthcare diagnostics. By leveraging the capabilities of deep learning in semantic understanding, especially in feature extraction and representation learning, this study introduces a new multimodal contrastive pretaining framework that aims to improve the quality and robustness of learned representations of 12-lead ECG signals. Our framework comprises two key components, including Cardio Query Assistant (CQA) and ECG Semantics Integrator(ESI). CQA integrates a retrieval-augmented generation (RAG) pipeline to leverage large language models (LLMs) and external medical knowledge to generate detailed textual descriptions of ECGs. The generated text is enriched with information about demographics and waveform patterns. ESI integrates both contrastive and captioning loss to pretrain ECG encoders for enhanced representations. We validate our approach through various downstream tasks, including arrhythmia detection and ECG-based subject identification. Our experimental results demonstrate substantial improvements over strong baselines in these tasks. These baselines encompass supervised and self-supervised learning methods, as well as prior multimodal pretraining approaches. Our work shows the potential of combining multimodal pretraining to improve the analysis of ECG signals.
@@ 1 Introduction
  • Compared to the strong baselines including the prior SOTA supervised and SSL methods, our evaluation demonstrates promising performances in arrhythmia detection and ECG-based user identification. For instance, we observe an improvement of 1.6% in AUC scores for diagnosing arrhythmia classes, and a 3.5% improvement in identifying subjects when compared to prior SSL methods.
@@ 2.1 Multimodal Representation Learning
Recent studies have introduced foundation models for integrating image and text, which involve both visual and vision-language pretraining. Models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), Florence (Yuan et al., 2021), and LiT (Zhai et al., 2022) demonstrate that dual-encoder architectures pretrained with contrastive objectives using image-text pairs, can develop robust representations for both modalities and improve the performances of image classification tasks. Beyond contrastive supervision, research involving encoder-decoder models trained with generative losses, such as (Yu et al., 2022b; Wang et al., 2021; 2022), has shown promising results in vision-language benchmarks, with the visual encoder remaining competitive in image classification tasks. Researchers have also applied these powerful representation learning techniques to medical imaging data, including radiography paired with clinical reports (Liu et al., 2023; You et al., 2023; Wan et al., 2024). However, compared to image-based applications, adapting multimodal pretraining methods for ECG signal processing remains relatively underexplored.
@@ 2.2.3 Multimodal Representation Learning in ECG
Although few, some studies have begun to explore the alignment of ECG signals with other modalities such as textual descriptions, EHR, and clinical notes (Li et al., 2023; Lalam et al., 2023; Liu et al., 2024). For instance, Lalam et al. (2023) utilized identical encoders to extract and contrastively align embeddings from ECG, EHR, and clinical notes. The pretrained model showed promising results in clinical diagnosis. Liu et al. (2024) adopted a similar approach to couple ECG signals with clinical notes and enhanced the ECG encoder’s effectiveness in zero-shot arrhythmia detection. However, the pretraining processes of these methods depend on costly annotations such as clinical notes and EHR, which are challenging to acquire on a large scale. Moreover, the variability in textual descriptions or reports associated with ECGs due to differences in detail, terminology, and style among clinicians and clinical contexts may complicate the learning of consistent mappings between ECG signals and text. The consequent variability could lead to misalignments between the ECG-text pairs. In this study, we propose to utilize a retrieval-augmented generation (RAG)-based pipeline to construct contextual ECG textual data without relying on costly notes and EHRs. Additionally, we introduce a captioning task in our model to achieve more nuanced representations.
@@ 3.1 Cardio Query Assistant (CQA) Framework
The Cardio Query Assistant (CQA) Framework, as shown in Figure 1, is designed to transform ECG condition labels into detailed descriptive text. The generated text incorporates demographic information, ECG conditions, and enriched waveform details. The developed CQA is outlined as follows:
@@ 3.1.2 Enhancement of ECG Semantics
The CQA Framework enriches the ECG-associated information through a comprehensive retrieval-augmented process. With the pre-constructed domain-knowledge database, the RAG-based approach enables CQA to query related knowledge using given information such as standard clinical labels, standard communications protocol for computer-assisted ECG (SCP) statements, diagnostic interpretations, and machine-generated reports associated with ECG data. For example, the potential lead-detailed waveform descriptions cannot be found in the SCP statements and arrhythmia diagnosis; however, the CQA framework queries this piece of enriched information based on the varying annotations and generates textual descriptions for the potential waveform. Further, this approach encourages the combinative use of ECG data from different sources without relying on labeled data, which means we can combine ECG databases with different annotations as a large-scale dataset for pretraining.
The output of the CQA framework is organized utilizing a Large Language Model (LLM), e.g., GPT-3.5, to generate the potential waveform details from the queried knowledge base. Take an example of the cardiac condition of the Right Bundle Branch Block (RBBB). By executing targeted queries in our database, it retrieves and generates descriptive context for specific waveform attributes. For example, the ECG-associated information of “RBBB” is queried and converted into related waveform features, including “prolonged QRS duration” and “M-shaped RSR’ pattern in leads V1-V3.”
@@ 3.2 Multimodal Contrastive Captioning with ECG Semantics Integrator (ESI) Framework
The ESI framework aims to improve the quality of representation extracted from ECG signals by pretraining a specialized ECG encoder alongside a textual encoder. This dual-modality training method has been proven by cutting-edge studies in contrastive language-image pretraining (CLIP) (Radford et al., 2021) and CoCa (Yu et al., 2022b) methodology.
For the ECG encoder, we have chosen a one-dimensional modified version of the ConvNext v2 architecture considering the sequential nature of ECG waveform data (Woo et al., 2023), which has been proven for its capacities of extracting both local and global contexts with the designed convolutional kernels. In parallel, the textual encoder utilizes BioLinkBERT, a derivative of the BERT architecture pretrained on biomedical texts, to effectively embed medical terminologies (Yasunaga et al., 2022).
@@ 3.2.1 Multimodal Contrastive Learning
Captioning Loss: While the dual-encoder approach encodes the text as an embedding for the contrastive learning purpose; the generative approach aims for detailed granularity and requires the model to predict the exact tokenized texts with ECG and preceding texts. This approach encourages the encoders to capture the semantic information embedded in the texts actively. Inspired by the image-text multimodal pretraining study CoCa (Yu et al., 2022b), we design to align the generated textual descriptions with the corresponding ECG signals by additionally defining a captioning loss $\mathcal{L}_{cap}$ similar to that used in image captioning tasks (Vinyals et al., 2015):
@@ 4 Evaluation
In this section, we describe our evaluation settings and experimental results. We first introduce the information on the datasets used and tasks performed in this study, along with the baseline methods we used in the comparisons. Our experiments explore three settings: zero-shot learning, linear probing (frozen features), and fine-tuning.
@@ 4.1 Training Setup
Pretraining Datasets. The proposed ESI signal encoder is pretrained from scratch. Therefore, the pretraining dataset directly impacts model’s generalizability. We constructed a large pretraining set combining three large-scale datasets with over 650,000 ECG-text training pairs. These datasets covers Chapman-Shaoxing (Zheng et al., 2020), PTB-XL (Wagner et al., 2020), and MIMIC-ECG (Gow et al., 2023). Each dataset contains 12-lead and 10-second ECG recordings sampled at 500 Hz. Here is a detailed breakdown of each dataset:
@@ 4.3.1 Fine-tuning & Linear Probing
Table 2 shows the performances of the evaluated methods. In the supervised learning methods, the ConvNeXt and ConvNeXt-tiny encoders show lower performance compared to specialized methods such as MLBF-Net and MVMSN. ConvNeXt-tiny outperforms the larger ConvNeXt model potentially due to overfitting with the smaller dataset. After pretraining, the ConvNeXt-based ESI method achieves the best performance on both PTB-XL and ICBEB datasets under both the frozen encoder and fine-tuning settings. Notably, the multimodal pretraining methods (ESI and MERL) significantly outperformed the signal-focused methods such as SimCLR and BYOL. This supports our hypothesis that signal-focused approaches may have limitations in learning robust and transferable representations for downstream tasks compared to the multimodal pretraining methods. The superior performance of the pretrained ESI encoder compared to the frozen encoder in supervised learning approaches also demonstrates the robustness of the learned features for arrhythmia diagnosis.
Table 2: Evaluation results of arrhythmia diagnosis task under different settings including supervised learning, linear probing (frozen encoder), and fine-tuning. The metric used is area under the ROC curve (AUC). The best results are highlighted in bold.
@@
LSTM (Strodthoff et al., 2020) 0.907 0.964
XResNet101 (Strodthoff et al., 2020) 0.925 0.974
ResNet50 (Strodthoff et al., 2020) 0.919 0.969
Ensemble (Strodthoff et al., 2020) 0.929 0.975
MLBF-Net (Zhang et al., 2021) 0.931 -
MVMSN (Yang et al., 2023) 0.933 -
ConvNeXt-Tiny (Woo et al., 2023) 0.917 0.972
ConvNeXt-Base (Woo et al., 2023) 0.913 0.969
@@
SimCLR (Chen et al., 2020) 0.764 0.788
BYOL (Grill et al., 2020) 0.771 0.804
CLOCS (Kiyasseh et al., 2021) 0.774 0.799
LEAVES (Yu et al., 2022a) 0.792 0.809
MERL (Liu et al., 2024) 0.887 -
(Ours) ESI-tiny 0.928 0.976
(Ours) ESI 0.932 0.977
@@
SimCLR (Chen et al., 2020) 0.919 0.971
BYOL (Grill et al., 2020) 0.924 0.969
CLOCS (Kiyasseh et al., 2021) 0.915 0.975
CRT (Zhang et al., 2022) 0.892 -
LEAVES (Yu et al., 2022a) 0.928 0.975
(Ours) ESI-tiny 0.936 0.978
(Ours) ESI 0.939 0.981
4.3.2 Zero-shot Inference
To further assess the learned representations during pretraining, we also evaluate the zero-shot learning inference assessment following METS (Li et al., 2023) and MERL (Liu et al., 2024). Other than the zero-shot evaluations, we include the few-shot setting of the existing signal-focused SSL approaches in the comparison, including SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), CLOCS (Kiyasseh et al., 2021), and LEAVES (Yu et al., 2022a), with 5% of the original training set as training samples on PTB-XL dataset.
Figure 3 demonstrates the zero-shot learning inference performance of the proposed ESI method alongside baselines (METS, MERL) and the few-shot fine-tuning results for signal-centered SSL methods. Our proposed method achieves the best performance on both AUC and macro F1 scores compared to all other methods. Additionally, ECG-text pretrained models generally outperformed signal-focused pretraining methods even without samples in fine-tuning. This highlights the improved robustness of representations from multimodal pretraining techniques.
Table 3: Evaluation results of arrhythmia diagnosis task under different settings in zero-shot learning on PTB-XL dataset. The metric used in this table is area under the ROC curve (AUC) and macro F1 score (F1-macro). $X-\%$ represents the percentage $X$ of the training set used as the training samples in fine-tuning the pretrained encoder. The best results are highlighted in bold.

### v2

@@ Abstract
The utilization of deep learning on electrocardiogram (ECG) analysis has brought the advanced accuracy and efficiency of cardiac healthcare diagnostics. In this work, we address a critical challenge in the field of ECG analysis with deep learning: learning robust representation without large-scale labeled datasets. We propose ECG Semantic Integrator (ESI), a novel multimodal contrastive pretraining framework that jointly learns from ECG signals and associated textual descriptions. ESI employs a dual objective function that comprises a contrastive loss and a captioning loss to develop representations of ECG data. To create a sufficiently large and diverse training dataset, we develop a retrieval-augmented generation (RAG)-based Large Language Model (LLM) pipeline, called Cardio Query Assistant (CQA). This pipeline is designed to generate detailed textual descriptions for ECGs from diverse databases. The generated text includes information about demographics and waveform patterns. This approach enables us to compile a large-scale multimodal dataset with over 660,000 ECG-text pairs for pretraining ESI, which then learns robust and generalizable representations of 12-lead ECG. We validate our approach through various downstream tasks, including arrhythmia detection and ECG-based subject identification. Our experimental results demonstrate substantial improvements over strong baselines in these tasks. These baselines encompass supervised and self-supervised learning methods, as well as prior multimodal pretraining approaches. Our work shows the potential of combining multimodal pretraining to improve the analysis of ECG signals.
The training code and generated waveform descriptions are available at https://github.com/comp-well-org/ESI.
@@ 1 Introduction
  • Compared to strong baselines including the prior SOTA supervised and SSL methods, our evaluation demonstrates promising performances in arrhythmia detection and ECG-based user identification. For instance, we observe an improvement of 1.6% in AUC scores for diagnosing arrhythmia classes, and a 3.5% improvement in identifying subjects when compared to prior SSL methods.
@@ 2.1 Multimodal Representation Learning
Recent studies have introduced foundation models for integrating image and text, which involve both visual and vision-language pretraining. Vision-language pretraining aims to learn multimodal foundation models with enhanced performance on downstream vision and language tasks. Models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), Florence (Yuan et al., 2021), and LiT (Zhai et al., 2022) use dual-encoder architectures where image-text pairs are mapped into a shared embedding space. These models are trained using a contrastive objective that pulls together embeddings of matched image-text pairs while pushing apart those of unmatched pairs. This approach has proven effective in developing robust and transferable representations for both image and text modalities, which significantly improves the performance on various vision-language tasks.
In parallel, other research has explored encoder-decoder architectures that leverage generative objectives. Models such as CoCa (Yu et al., 2022b), SimVLM (Wang et al., 2021), BLIP (Li et al., 2022), BLIP-2 (Li et al., 2023b) and OFA (Wang et al., 2022) integrate both encoding and decoding processes, which enables the model to generate text from visual inputs. These models use language modeling alongside image-text matching and generate semantically meaningful outputs. This line of study also demonstrates a flexible and powerful potential for multimodal learning and demonstrates strong performance across a variety of vision-language benchmarks.
The power of these multimodal pretraining techniques has not been limited to general image and text data. Researchers have adapted these approaches to medical imaging, particularly in radiography paired with clinical reports (Liu et al., 2023; You et al., 2023; Wan et al., 2024). However, despite the progress in image-based applications, applying these multimodal pretraining methods to ECG signal processing remains relatively underexplored. In this study, we extend these advanced techniques to ECG data for more robust and generalizable representations in clinical settings.
@@ 2.2.3 Multimodal Representation Learning in ECG
Although few, some studies have begun to explore the alignment of ECG signals with other modalities such as textual descriptions, EHR, and clinical notes (Li et al., 2023a; Lalam et al., 2023; Liu et al., 2024). For instance, Lalam et al. (2023) utilized identical encoders to extract and contrastively align embeddings from ECG, EHR, and clinical notes. The pretrained model showed promising results in clinical diagnosis. Liu et al. (2024) adopted a similar approach to couple ECG signals with clinical notes and enhanced the ECG encoder’s effectiveness in zero-shot arrhythmia detection. However, the pretraining processes of these methods depend on costly annotations such as clinical notes and EHR, which are challenging to acquire on a large scale. Moreover, the variability in textual descriptions or reports associated with ECGs due to differences in detail, terminology, and style among clinicians and clinical contexts may complicate the learning of consistent mappings between ECG signals and text. The consequent variability could lead to misalignments between the ECG-text pairs. In this study, we propose to utilize a retrieval-augmented generation (RAG)-based pipeline to construct contextual ECG textual data without relying on costly notes and EHRs. Additionally, we introduce a captioning task in our model to achieve more nuanced representations.
@@ 3 Methods
Figure 2: Example of a 12-lead ECG signal and its associated metadata. The left side displays the 12-lead ECG waveform, which provides a visual representation of the heart’s electrical activity. The right side includes relevant patient demographic information (age, gender), a summary of ECG metadata including high-level clinical findings (e.g., Sinus rhythm, Abnormal R-wave progression), and a detailed textual description of the ECG conditions generated with the CQA pipeline.
@@ 3.1 Cardio Query Assistant (CQA) Framework
The Cardio Query Assistant (CQA) Framework, as shown in Figure 1, is designed to transform ECG condition labels into detailed descriptive text. The generated text incorporates demographic information, ECG conditions, and enriched waveform details. An example of a raw ECG signal, its associated metadata, and the generated ECG waveform description is provided in Figure 2. The developed CQA is outlined as follows:
@@ 3.1.2 Enhancement of ECG Semantics
The CQA Framework enriches the ECG-associated information through a comprehensive retrieval-augmented process. With the pre-constructed domain-knowledge database, the RAG-based approach enables CQA to query related knowledge using given information such as standard clinical labels, standard communications protocol for computer-assisted ECG (SCP) statements, diagnostic interpretations, and machine-generated reports associated with ECG data. SCP statements are standardized textual formats that provide consistent documentation of ECG findings, following international protocols for computer-assisted ECG interpretation. While these datasets provide valuable diagnostic information, they often lack explicit details about the waveform patterns that are critical for a thorough ECG analysis.
To address this gap, the CQA Framework employs a RAG approach that enables it to query the knowledge database, which is introduced in the prior section, for relevant information and generate comprehensive textual descriptions of the potential waveform characteristics. For instance, SCP statements and standard arrhythmia diagnoses may not directly include detailed waveform descriptions; however, by querying the domain-knowledge database, the CQA Framework can synthesize this enriched information, producing detailed descriptions of the waveform patterns that correspond to specific ECG conditions.
With the constructed RAG pipeline in CQA, we asked the waveform information by using simple prompts of "How is ECG Condition reflected in a 12-lead ECG"? The output of the CQA framework is organized utilizing an LLM, e.g., GPT-3.5, to generate the potential waveform details from the queried knowledge base. Take an example of the cardiac condition of the Right Bundle Branch Block (RBBB). By executing targeted queries in our database, it retrieves and generates descriptive context for specific waveform attributes. For example, the ECG-associated information of “RBBB” is queried and converted into related waveform features, including “prolonged QRS duration” and “M-shaped RSR’ pattern in leads V1-V3.”
@@ 3.2 Multimodal Contrastive Captioning with ECG Semantics Integrator (ESI) Framework
The ESI framework aims to improve the quality of representation extracted from ECG signals by pretraining a specialized ECG encoder alongside a textual encoder. This dual-modality training method has been proven by cutting-edge studies in contrastive language-image pretraining (CLIP) (Radford et al., 2021) and CoCa (Yu et al., 2022b) methodology. Inspired by these two studies, we developed our pretraining architecture with both contrastive and generative objectives. While we recognize there are newer approaches in multimodal contrastive learning beyond CLIP and CoCa as discussed in Section 2.1, the choice of development can be motivated by several factors. First, the CLIP structure is well-proven by multiple prior works, including pretraining language with medical images (Liu et al., 2023; You et al., 2023; Wan et al., 2024) and physiological signals (Li et al., 2023a; Lalam et al., 2023; Liu et al., 2024). The core contrastive objective of CLIP is well-suited for our primary goal of aligning ECG signals with their textual interpretations. Moreover, the relative simplicity of the CLIP architecture allows for seamless integration of our captioning loss inspired by CoCa (Yu et al., 2022b). We further validate the contribution of the contrastive and generative objective via ablation study as shown in Section 5.1.
For the ECG encoder, we have chosen a one-dimensional modified version of the ConvNext v2 architecture considering the sequential nature of ECG waveform data (Woo et al., 2023), which has been proven for its capacities of extracting both local and global contexts with the designed convolutional kernels. In parallel, the textual encoder utilizes BioLinkBERT, a derivative of the BERT architecture pretrained on biomedical texts, to effectively embed medical terminologies (Yasunaga et al., 2022). BioLinkBERT is an extension of the standard BERT model (Devlin et al., 2018), specifically designed to improve the understanding of biomedical texts. Unlike traditional BERT, which processes each document independently, BioLinkBERT is pretrained on biomedical literature from PubMed, which takes advantage of the natural links between documents such as citations and references. This pretraining strategy motivates us to leverage BioLinkBert for tasks that require a deep understanding of biomedical concepts and terminology.
@@ 3.2.1 Multimodal Contrastive Learning
Captioning Loss: While the dual-encoder approach encodes the text as an embedding for the contrastive learning purpose, the generative approach aims for detailed granularity and requires the model to predict the exact tokenized texts with ECG and preceding texts. This approach encourages the encoders to capture the semantic information embedded in the texts actively. Inspired by the image-text multimodal pretraining study CoCa (Yu et al., 2022b), we design to align the generated textual descriptions with the corresponding ECG signals by additionally defining a captioning loss $\mathcal{L}_{cap}$ similar to that used in image captioning tasks (Vinyals et al., 2015):
@@ 4 Evaluation
In this section, we describe our evaluation settings and experimental results. We first introduce the information on the datasets used and tasks performed in this study, along with the baseline methods we used in the comparisons. Our experiments explore three settings: zero-shot learning, linear probing (frozen features), and fine-tuning. In our experiments, we conduct multiple resampling runs to assess the robustness of the model, particularly for the linear probing and fine-tuning settings. Given that the encoder is frozen during linear probing evaluations, the randomness is only applied to the output linear heads during the linear probing setup. For each setting, we average the results over five different runs and report the standard deviations alongside the mean performance metrics.
@@ 4.1 Training Setup
Pretraining Datasets. The proposed ESI signal encoder is pretrained from scratch. Therefore, the pretraining dataset directly impacts the model’s generalizability. We constructed a large pretraining set combining three large-scale datasets with over 650,000 ECG-text training pairs. These datasets covers Chapman-Shaoxing (Zheng et al., 2020), PTB-XL (Wagner et al., 2020), and MIMIC-ECG (Gow et al., 2023). Each dataset contains 12-lead and 10-second ECG recordings sampled at 500 Hz. Here is a detailed breakdown of each dataset:
@@ 4.3.1 Fine-tuning & Linear Probing
Table 2 shows the performances of the evaluated methods. In the supervised learning methods, the ConvNeXt and ConvNeXt-tiny encoders show lower performance compared to specialized methods such as MLBF-Net and MVMSN. ConvNeXt-tiny outperforms the larger ConvNeXt model, potentially due to overfitting with the smaller dataset. After pretraining, the ConvNeXt-based ESI method achieves the best performance on both PTB-XL and ICBEB datasets under both the frozen encoder and fine-tuning settings. Notably, the multimodal pretraining methods (ESI and MERL) significantly outperformed the signal-focused methods such as SimCLR and BYOL. This supports our hypothesis that signal-focused approaches may have limitations in learning robust and transferable representations for downstream tasks compared to the multimodal pretraining methods. The superior performance of the pretrained ESI encoder compared to the frozen encoder in supervised learning approaches also demonstrates the robustness of the learned features for arrhythmia diagnosis.
4.3.2 Zero-shot Inference
To further assess the learned representations during pretraining, we also evaluate the zero-shot learning inference assessment following METS (Li et al., 2023a) and MERL (Liu et al., 2024). Other than the zero-shot evaluations, we include the few-shot setting of the existing signal-focused SSL approaches in the comparison, including SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), CLOCS (Kiyasseh et al., 2021), and LEAVES (Yu et al., 2022a), with 5% of the original training set as training samples on PTB-XL dataset.
Table 3 demonstrates the zero-shot learning inference performance of the proposed ESI method alongside baselines (METS, MERL) and the few-shot fine-tuning results for signal-centered SSL methods. Our proposed method achieves the best performance on both AUC and macro F1 scores compared to all other methods. Additionally, ECG-text pretrained models generally outperformed signal-focused pretraining methods even without samples in fine-tuning. This highlights the improved robustness of representations from multimodal pretraining techniques.
Table 2: Evaluation results of arrhythmia diagnosis task under different settings including supervised learning, linear probing (frozen encoder), and fine-tuning. The metric used is the area under the ROC curve (AUC). The best results are highlighted in bold.
@@ 4.3.2 Zero-shot Inference
LSTM (Strodthoff et al., 2020) 0.927±0.013 0.964±0.015
XResNet101 (Strodthoff et al., 2020) 0.928±0.013 0.974±0.013
ResNet50 (Strodthoff et al., 2020) 0.930±0.015 0.969±0.015
Ensemble (Strodthoff et al., 2020) 0.934±0.013 0.975±0.013
MLBF-Net (Zhang et al., 2021) 0.931±0.021 -
MVMSN (Yang et al., 2023) $-{}_{\pm\text{-}}$
ConvNeXt-Tiny (Woo et al., 2023) 0.918±0.016 0.970±0.012
ConvNeXt-Base (Woo et al., 2023) 0.914±0.014 0.970±0.012
@@
SimCLR (Chen et al., 2020) 0.766±0.016 0.791±0.012
BYOL (Grill et al., 2020) 0.776±0.014 0.800±0.014
CLOCS (Kiyasseh et al., 2021) 0.777±0.012 0.802±0.015
LEAVES (Yu et al., 2022a) 0.792±0.015 0.809±0.013
MERL (Liu et al., 2024) $-{}_{\pm\text{-}}$
(Ours) ESI-tiny 0.927±0.009 0.975±0.006
(Ours) ESI 0.931±0.008 0.978±0.006
@@
SimCLR (Chen et al., 2020) 0.916±0.015 0.968±0.016
BYOL (Grill et al., 2020) 0.925±0.014 0.971±0.014
CLOCS (Kiyasseh et al., 2021) 0.918±0.013 0.977±0.012
CRT (Zhang et al., 2022) $-{}_{\pm\text{-}}$
LEAVES (Yu et al., 2022a) 0.926±0.012 0.976±0.013
(Ours) ESI-tiny 0.935±0.011 0.978±0.010
(Ours) ESI 0.938±0.011 0.982±0.009
Table 3: Evaluation results of arrhythmia diagnosis task under different settings in zero-shot learning on PTB-XL dataset. The metric used in this table is the area under the ROC curve (AUC) and macro F1 score (F1-macro). $X-\%$ represents the percentage $X$ of the training set used as the training samples in fine-tuning the pretrained encoder. The best results are highlighted in bold.